OpenStudy (anonymous):

a-2/63 = 1/a : solve for a

OpenStudy (anonymous):

\[\frac{a-2}{63}=\frac{1}{a}\] Use cross multiplication: \[\frac{X}{Y}=\frac{C}{D}\] XD = CY